### v2

# ICD-LM: Configuring Vision-Language In-Context Demonstrations by Language Modeling

Abstract
This paper studies how to configure powerful In-Context Demonstration (ICD) sequences for a Large Vision-Language Model (LVLM) to solve Vision-Language tasks through In-Context Learning (ICL). After observing that configuring an ICD sequence is a mirror process of composing a sentence, i.e., just as a sentence can be composed word by word via a Language Model, an ICD sequence can also be configured one by one. Consequently, we introduce an ICD Language Model (ICD-LM) specifically designed to generate effective ICD sequences. This involves creating a dataset of hand-crafted ICD sequences for various query samples and using it to train the ICD-LM. Our approach, diverging from traditional methods in NLP that select and order ICDs separately, enables to simultaneously learn how to select and order ICDs, enhancing the effect of the sequences. Moreover, during data construction, we use the LVLM intended for ICL implementation to validate the strength of each ICD sequence, resulting in a model-specific dataset and the ICD-LM trained by this dataset is also model-specific. We validate our methodology through experiments in Visual Question Answering and Image Captioning, confirming the viability of using a Language Model for ICD configuration. Our comprehensive ablation studies further explore the impact of various dataset construction and ICD-LM development settings on the outcomes. The code is given in https://github.com/ForJadeForest/ICD-LM.
1 Introduction
With the escalation in model size [4, 45, 7, 6, 32, 44], the capabilities of Large Language Models (LLMs) have been largely enhanced. One salient emerging ability is In-Context Learning (ICL) [38, 39, 8], which can make LLMs quickly adapt to novel tasks using a few examples that are termed In-Context Demonstrations (ICDs). Crucially, ICL does not need gradient updates to model parameters, thereby circumventing challenges tied to additional data collection and model fine-tuning. Inspired by such success in NLP, researchers in the Vision-Language (VL) domain also design Large Vision-Language Models (LVLMs) capable of ICL [47, 33, 1, 2], paving the way for novel research and applications. Then an elemental question arising from these trained LVLMs is how to achieve optimal ICL performance across various vision-language tasks.
Figure 1: (a) The traditional ICD configuration methods separately select and order the ICDs, leading to sub-optimal ICL performance. (b) Our ICD-LM enables the step-by-step generation of ICD configurations and simultaneously considers the selection of ICDs and the ordering of ICD sequences.
Before delving into ICL performance optimization in VL, we first retrospect the studies on ICL in NLP. While well-trained LLMs are identified as good In-Context Learners [4, 21], it has been observed that diverse ICD configurations heavily affect the ICL performance. For example, using different examples from the supporting set as ICDs [18, 9, 23], or even changing the orders of these ICDs [20, 48, 11], may lead to significant performance fluctuations. Driven by these findings, researchers design diverse heuristic-based (e.g., embedding similarity-based) retrieval methods [18, 28, 30] to select representative examples as the ICDs. Then these ICDs are reordered based on some principles, e.g., the Minimal Description Length [40] or Global and Local Entropy [20], for further improving the results.
While employing these methods to configure the ICDs yields consistent improvements, they still suffer from a few limitations. First, the independent operation of retrieval and reordering processes potentially yields sub-optimal results. For example, using similarity-based retrieval methods followed by reordering of ICD configuration may lead to the sub-optimal ICL performance, as shown in Fig. 1 (a) (i), where colors closer to the query sample (grey) in the figure indicate higher similarity. Second, as shown in Fig. 1 (a) (ii), researchers have noted that different LLMs exhibit sensitivities to particular ICD configurations [20], suggesting that the ICD configuration should be model-specific. Yet, most retrieval or reordering strategies remain model-agnostic, potentially leading to sub-optimal results.
Then if we want to directly extend these NLP strategies to VL, the above-mentioned challenges will be further exacerbated by the multimodal combinatorial complexity of vision and language data. For instance, although retrieving similar text samples typically yields favorable results in NLP, simply retrieving similar single-modal data, such as images, may not guarantee the best performance in VL tasks [41]. However, attempting to match both visual and textual modalities during retrieval will further magnify the time complexity. Additionally, VL tasks have more diverse problem structures than NLP, necessitating distinct retrieval or reordering strategies tailored to each VL task, thereby elevating the design and labor overhead.
In this study, we address these issues by introducing a novel approach for configuring ICDs to solve diverse VL tasks. Before delving into our approach, we reconsider the ideal process of constructing an ICD sequence. Given the query sample, we hope to subsequently select a series of ICDs that can maximize the ICL performance. This iterative selection mirrors the construction of a sentence wherein the word is selected one by one to ensure fluency. Recognizing this analogy, we propose to treat the ICD configuration as a specialized Language Model (LM). The only difference is that the vocabulary of this ICD-LM is not conventional words, but are the examples from the supporting set that will be used as ICDs.
We show the ICD-LM pipeline in Fig. 1 (b). To deploy this ICD-LM, a primary challenge arises that we do not have a ”ground-truth” dataset that indicates which examples or their orders can form an optimal ICD sequence for a specific LVLM-based in-context learner. To address it, we construct a dataset containing the pairs of a query and the corresponding ICD sequence. During construction, we use this frozen specific LVLM to measure whether an ICD sequence is good for the query or not. In this way, the constructed dataset becomes model-specific. Then we use this dataset to train a Transformer-based ICD-LM where each token in the vocabulary represents an example from the supporting set. During training, the ICD-LM concurrently learns the selection and ordering of ICDs, eliminating the necessity to operate two separate stages for constructing an ICD sequence as previous methods.
Besides alleviating the above-mentioned limitations, we also find some other interesting characteristics of our method. First, we evaluate the length extrapolation ability of ICD-LM, which pertains to its performance when generating ICD sequences that exceed the maximum length represented in the dataset. This is contrasted with its interpolation ability,i.e., the performance within the maximum ICD length of that dataset. For instance, if a dataset only contains 2-shot ICDs, the interpolation ability denotes the effect of the 1- and 2-shot ICDs when implementing ICL, while the extrapolation ability denotes the effect of more shot ICDs like 4- or even 8-shot ICDs. Interestingly, we find that the trained ICD-LM has strong length extrapolation ability, e.g., when trained on a dataset containing only 2-shot ICDs, it can proficiently configure 4-shot ICDs. Second, the trained ICD-LM may configure a fixed ICD sequence for each query while achieving excellent ICL performance. For example, in the IC/VQA tasks, the golden ICD sequence achieves 6.91/1.24 improvements compared with a strong baseline that retrieves the ICDs by similar images.
2 Related Work
Models with In-Context Learning Ability. Prompt engineering enables pre-trained Large Language Models(LLMs) to address downstream tasks without the need for fine-tuning [24, 4, 29]. A variant, ICL, enhances this ability by constructing prompts with a few examples. This has been demonstrated in LLMs such as GPT-3 [4], LLaMA [32], and MPT [31]. Recently, witnessing such success in NLP, the VL domain has also developed numerous LVLMs with prompt engineering abilities [49, 17, 13, 1, 33, 43]. With the support of LLM, Frozen [33] and Flamingo [1] have developed the ability for ICL. Among these LVLMs, Flamingo simulates the ICL approach during training, resulting in a more powerful ICL ability. However, due to the unavailability of an official open-source version of Flamingo, we use an unofficial implementation, OpenFlamingo [2] in this study to explore whether ICD-LM can configure useful ICD sequences.
Configuring In-Context Demonstrations. Although ICL assists LLMs in better adapting to downstream tasks, its performance is highly sensitive to the selection [18, 9, 23] and ordering [20, 48, 11] of ICDs. Numerous studies have explored diverse methods to select ICDs in the NLP field. For example,  [18] selects ICDs based on the embedding similarity between ICDs and test samples where the embeddings are extracted from an existent language encoder. Such method is further developed by training an encoder specifically for selection [26, 15, 42, 36, 15, 37, 5]. Additionally,  [40, 22, 14] leverage LLM outputs to select and  [46] formulates ICD selection as a Markov decision process [3].
Regarding the ordering of ICDs, researchers propose some principles to measure the quality of ICD configurations, e.g., the Minimal Description Length [40] and Global and Local Entropy [20]. However, these reordering methods are limited to classification tasks in NLP, as they require the calculation of conditional probabilities for each label given the task inputs. These approaches are inapplicable to certain tasks lacking a predefined label space, such as open-ended question answering or text generation.
Besides these NLP studies, in VL, [41] explores diverse ICD configurations on image captioning, while only the heuristic-based methods are used for selecting ICDs and do not consider the ordering. In contrast, our ICD-LM can simultaneously learn how to select and reorder the samples and moreover, our ICD-LM is model-specific.
3 Model-Specific ICD-LM
In this section, we introduce how to build a model-specific ICD-LM for configuring the ICD sequence for a given LVLM. First, we briefly introduce the formulations of ICL for Vision-Language (VL) tasks in section 3.1. Then we show how to construct a model-specific dataset (section 3.2) that will be used to train the ICD-LM, whose structure will be detailed in Section 3.3.
3.1 In-Context Learning (ICL)
Given a query input $\mathbf{x}^{\prime}$, ICL predicts the corresponding output $\mathbf{y}^{\prime}$ using a well-trained foundation model $\mathcal{M}$, conditioned on the concatenation of an in-context sequence $\mathcal{S}$ and this query. If the in-context sequence contains $K$-shot ICDs, it is denoted it as: $\mathcal{S}^{K}=\{\hat{\bm{d}}_{1},\hat{\bm{d}}_{2},...,\hat{\bm{d}}_{K}\}$, where $\hat{\bm{d}}_{k}$ is the $k$-th ICD. Then ICL can be formulated as:
$\mathbf{y}^{\prime}\leftarrow\bm{P}_{\mathcal{M}}(\mathbf{y}^{\prime}\mid\mathcal{S}^{k},\mathbf{x}^{\prime}),$ (1)
where $\bm{P}_{\mathcal{M}}$ denotes the predicted probability of $\mathcal{M}$ and $\leftarrow$” represents the decoding strategy, e.g., beam search.
For each ICD $\hat{\bm{d}}$, it is selected from a supporting dataset $\mathcal{D}_{S}=\{\bm{d}_{1},...,\bm{d}_{N}\}$, where each sample $\bm{d}_{i}=(\mathbf{x}_{i},\mathbf{y}_{i})$: $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$ respectively denote the input and the corresponding label. It is noteworthy that in diverse VL tasks, $\mathbf{x}$ and $\mathbf{y}$ have different forms. For instance, in Image Captioning (IC), $\mathbf{x}$ is the image and $\mathbf{y}$ is the caption; and in Vision Question Answering (VQA), $\mathbf{x}$ contains the image and the question, while $\mathbf{y}$ is the answer.
3.2 Constructing the Model-Specific Dataset
Figure 2: The pipeline of constructing $\mathcal{D}_{\mathcal{M}}$. The darker color of $S^{K}_{i,j}$ indicates a higher score given by Eq. 2 in figure (c).
To train our model-specific ICD-LM for generating effective ICD sequences for a given LVLM $\mathcal{M}$, we should first construct a model-specific dataset $\mathcal{D}_{\mathcal{M}}$ containing high-quality ICD sequences for different query inputs. To avoid confusion, Figure 2 shows how to construct $\mathcal{D}_{\mathcal{M}}$ from an VL dataset, e.g., the COCO dataset [16] for IC. Formally, given an dataset $\mathcal{D}$ which is already split into the training part $\mathcal{D}_{R}$ and the test part $\mathcal{D}_{E}$, we build $\mathcal{D}_{\mathcal{M}}$ only from the training part $\mathcal{D}_{R}$.
As shown in Fig. 2 (a), initially, we randomly select $n$ samples from $\mathcal{D}_{R}$ to form an anchor set $\mathcal{A}$. Then for each sample $\bm{a}_{m}=\{\bm{x}_{m},\bm{y}_{m}\}\in\mathcal{A}$, we aim to construct a $K$-shot in-context sequence $\mathcal{S}_{m}^{K}$ for it. Then $\mathcal{D}_{\mathcal{M}}=\{(\bm{a}_{1},\mathcal{S}_{1}^{K}),(\bm{a}_{2},\mathcal{S}_{2}^{K}),...,(\bm{a}_{M},\mathcal{S}_{M}^{K})\}$ where each training sample contains an query $\bm{a}_{m}$ and the corresponding $K$-shot in-context sequence $\mathcal{S}_{m}^{K}$.
To avoid confusion, we remove the subscript $m$ in following. To construct $\mathcal{S}^{K}=\{\bm{d}_{1},...,\bm{d}_{K}\}$, we need to select $K$-shot samples from the supporting set $\mathcal{D}_{S}$, which is set to the complement set of $\mathcal{A}$ in $\mathcal{D}_{R}$: $\mathcal{D}_{R}\backslash\mathcal{A}$. Meantime, we also need to decide which samples should be selected in turn. To achieve this, given the anchor sample $\bm{a}=\{\bm{x},\bm{y}\}$ and the partially constructed in-context sequence, e.g., a $k-1$-shot $\mathcal{S}^{k-1}$, we need to know that after adding which sample $\bm{d}\in\mathcal{D}_{S}$, the ICL performance improvement can be maximized by applying the given LVLM $\mathcal{M}$:
$\small\hat{\bm{d}}_{k}=\mathop{\arg\max}\limits_{\bm{d}\in\mathcal{D}_{S}}I_{\mathcal{M}}(\{\bm{d},\mathcal{S}^{k-1}\},\bm{a})-I_{\mathcal{M}}(\mathcal{S}^{k-1},\bm{a}),$ (2)
where $I_{\mathcal{M}}$ is one kind of ICL performance measurement related to $\mathcal{M}$. Note that Eq. (2) actually uses the greedy sampling method to select the samples every time, while we can use beam search here to further achieve a better solution. Additionally, to improve the diversity of the dataset, we will keep the top-$b$ highest-scoring ICD sequences $\{S^{K}_{m,1},S^{K}_{m,2},...,S^{K}_{m,b}\}$ at the last iteration for each $a_{m}$, where $b$ is equal to the beam size. For example, when setting beam size to $5$ as shown in Fig. 2 (c), we can get $5$ diverse high-quality ICD sequences for an anchor sample.
Intuitively, for diverse tasks, we can use the corresponding “golden measurement” as $I_{\mathcal{M}}$, e.g., setting it to CIDEr [35] for IC and accuracy for VQA. However, this strategy encounters two limitations. The first one is that for diverse VL tasks, we need diverse corresponding measurements, which is inconvenient. Second, some “golden measurements” may be impractical to deploy. For example, for IC, calculating CIDEr requires the LVLM to forward multiple times to sample an integral sentence, and then it costs expensive time burdens to construct the dataset. While for VQA, accuracy is a binary value (accuracy=1 when correct and 0 when wrong), then maybe lots of candidates in $\mathcal{D}_{S}$ will make the accuracy change from 0 to 1 and then it is hard to judge which one of them is the most suitable one.
To overcome these two limitations, we use a relatively general measurement as $I_{\mathcal{M}}$. Formally, since we have the ground-truth results $\bm{y}$ of the anchor sample, we can use the given LVLM $\mathcal{M}$ to measure the prediction confidence of $\bm{y}$ given the input $\bm{x}$ and the in-context sequence $\mathcal{S}^{K}$:
$I_{\mathcal{M}}(\mathcal{S}^{K},\bm{a})$ $=P_{\mathcal{M}}(\bm{y}|\mathcal{S}^{K},\bm{x})$ (3)
$=\prod_{t}P_{\mathcal{M}}(y^{(t)}|\mathcal{S}^{K},\bm{x},y^{(1:t-1)}).$
Note that in VL tasks, the ground-truth label $\bm{y}=\{y^{(1)},...,y^{(T)}\}$ is usually a text sequence, thus we can decompose the probability distribution into a series of productions. In this way, Eq. (2) selects a sample that can further maximize the prediction confidence given the query input and the current in-context sequence.
In implementation, $\mathcal{D}_{S}$ usually contains huge amounts of samples, e.g., $\mathcal{D}_{S}$ in COCO [16] contains about $10^{5}$ samples. However, we need to calculate Eq. (2) for each $\bm{d}\in\mathcal{D}_{S}$ when selecting $\hat{\bm{d}}_{k}$ for each $\bm{a}\in\mathcal{A}$, which means the whole process of building $\bm{D}_{\mathcal{M}}$ is quite time-consuming. To alleviate the cost, as shown in Fig. 2 (b), for each specific $\bm{a}$, we narrow the set size by sampling a much smaller subset $\mathcal{D}_{S}^{\bm{a}}$, e.g., containing 64 samples $\mathcal{D}_{S}^{\bm{a}}=\{d_{\bm{a}}^{1},d_{\bm{a}}^{2},...,d_{\bm{a}}^{64}\}$, from $\mathcal{D}_{S}$ for selecting $\hat{\bm{d}}_{k}$. We use diverse sampling strategies to construct this subset, e.g., retrieving some similar samples as $\bm{a}$, and implement exhaustive ablation studies to explore which strategies are useful in Section 4.4.
3.3 Training Model-Specific ICD-LM
Figure 3: Top: The architecture of ICD-LM, which is a two-layer Transformer decoder blocks. Bottom: The components of the input embeddings, which is the sum of the random initialized learnable embeddings, the image and text embeddings extracted by CLIP. The dotted block means that some tasks do not exist the text input, e.g., IC.
After constructing the model-specific dataset $\mathcal{D}_{\mathcal{M}}$, we use it to train our ICD-LM, as Figure 3 shows, it is a two-layer Transformer decoder blocks [34]. The primary difference between our ICD-LM and the traditional LM lies in the tokens of the vocabulary. Specifically, the tokens in our vocabulary are the samples from the supporting set $\mathcal{D}_{S}$, e.g., the first token corresponds to the first sample in $\mathcal{D}_{S}$.
[…]
To train the ICD-LM, we first need to embed the tokens of the vocabulary to get dense embeddings. Since each token contains both image and text, we use the vision encoder $F_{I}(\cdot)$ and the language encoder $F_{T}(\cdot)$ of CLIP [25] to embed the image and text, respectively. Meanwhile, we add each of these embeddings with a learnable part which is randomly initialized. Then for the $i$-th token $\bm{d}_{i}=(I_{i},T_{i})$ in the vocabulary where $I_{i}$ and $T_{i}$ are the corresponding image and text, respectively, its token embedding is $\bm{e}_{i}$:
$\bm{e}_{i}=F_{I}(I_{i})+F_{T}(T_{i})+\bm{r}_{i},$ (4)
where $\bm{r}_{i}$ is the randomly initialized learnable embedding for the $i$-th token. Note that $T_{i}$ varies between IC and VQA tasks. In IC, $T_{i}$ represents the caption, whereas in VQA, it denotes the question and the answer.
For the test query $x^{\prime}$, we also use the same vision and language encoders to embed it. For VQA, the image and question will be embedded and summed, while for IC, only the image is embedded. Lastly, we employ the cross-entropy loss for training as a standard LM that given the previously $k-1$ tokens, we maximize the probability of the $k$-th ground-truth token.
3.4 Configuring the ICD Sequence
[…]
4 Experiments
4.1 Tasks and Datasets
Our approach is evaluated on MS-COCO for Image Captioning (IC) and VQAV2 for Visual Question Answering (VQA). For each corresponding dataset, we use the train split to construct the $\mathcal{D}_{\mathcal{M}}$ for training ICD-LM and use the validation split to evaluate the performance of ICD configurations generated by ICD-LM.
MS-COCO [16]: MS-COCO is widely used in IC, which is divided into $118,287$ training, $5,000$ validation, and $5,000$ test image-caption pairs. Notably, each training image is associated with five distinct human-annotated captions.
VQAV2 [10]: VQAV2 emphasizes open-ended VQA tasks, which encompasses $4,437,570$ question-answer pairs in its training split, supplemented by an additional $2,143,540$ pairs in the validation split.
4.2 Implementation Details
Constructing $\mathcal{D}_{\mathcal{M}}$: We construct 2-shot $\mathcal{D}_{\mathcal{M}}$ for main experiments. We select $5000$ samples to get the anchor set $\mathcal{A}$. For each anchor sample, we randomly choose $64$ samples to build the sub-supporting set $\mathcal{D}^{\bm{a}}_{S}$. The beam size for these processes is 5. More details are in Appendix.
Training ICD-LM: Different training strategies are employed for IC and VQA when training ICD-LM. In IC, the weight of CLIP model will be frozen, and an MLP adapter is introduced to its output. While, for VQA, the CLIP encoder remains trainable, and no adapter is appended. Both the ICD-LM for VQA and IC use 2-layer Transformer decoder blocks [34] or LSTM [27] as the architecture of LM. The training phase leverages the AdamW optimizer [19] and a cosine learning rate scheduler. We set the learning rate to $1\times 10^{-4}$ and the batch size to 128. We train our ICD-LM for 20 epochs. All training processes are carried out with mixed precision and 2 RTX3090 GPUs.
Implementing ICL: To implement ICL, we use OpenFlamingoV2-9B [2] as our LVLM. We use beam search during inference where the beam size is set to 3. Besides, we set the maximum number of generated tokens as 20 in IC and 5 in VQA. All experiments are deployed on an RTX 3090, operating in BF16 mode. The prompt structures of IC and VQA are shown in Table  1.
Table 1: Prompt Formats for IC and VQA tasks with placeholders.
Task Prompt Format
VQA <image>Question:<Q> Short answer:<A>
Caption <image>Output:<X>
4.3 Results and Analyses
4.3.1 Comparison Methods
We compare ICD-LM to 4 ICD selection strategies:
Random Sample (RS): RS constructs $\mathcal{S}^{k}$ by randomly selecting and ordering $k$ ICDs from $\mathcal{D}_{S}$.
Similarity-based Retrieval methods: These methods form $\mathcal{S}^{k}$ by computing the similarity between the query input $\mathbf{x}^{\prime}$ and ICDs in $\mathcal{D}_{S}$. Specifically, we use CLIP to extract the features of text or image and use cosine similarity to measure the similarity between the extracted features. We follow  [18], sorting examples in ascending order by their similarity to the query input, so the rightmost demonstration is the closest example. (1). Similarity-based Image-Image Retrieval (SIIR): We select $k$ ICDs from $\mathcal{D}_{S}$ with highest image similarity to the query image. (2). Similarity-based Text-Text Retrieval (STTR): We select $k$ ICDs from $\mathcal{D}_{S}$ that have most similar texts to the text of the query. For VQA, with question as input, we can use STTR to form $\mathcal{S}^{k}$. While for IC, the query input is only an image without any text input, thus STTR is infeasible. (3). Similarity-based Image-Text Retrieval (SITR): Taking advantage of the cross-modal retrieval capability of CLIP, we can compute the similarity between different modalities. Specifically, we compute the similarity between query image and all text of $d_{i}\in\mathcal{D}_{S}$ and select ICDs whose texts have the top-$k$ similarities with the query image. The text input varies per task, e.g., it is the caption for IC and the question for VQA.
4.3.2 Main Result
[…]
Overall, ICD-LM achieves the best performance on most cases compared to other methods. Notably, ICD-LM consistently outperforms other methods for the Avg:1$\sim$2. Specifically, in IC, ICD-LM surpasses the best similarity-based retrieval method, SIIR, by 6.03 in CIDEr score (84.32 vs. 78.29). Similarly, for VQA, ICD-LM exceeds the performance of the best similarity-based retrieval method, STTR, by achieving an accuracy 3.07 higher (48.75 vs. 45.68). The results indicate that ICD-LM performs well in generating 2-shot ICD configurations.
Moreover, we find that ICD-LM has remarkable extrapolation abilities. Regarding the Avg:3$\sim$8 metric, ICD-LM maintains the top performance in both IC and VQA. Specially, compared with SIIR, ICD-LM achieves 0.8 higher CIDEr in IC (96.52 vs. 95.72) and 0.75 higher accuracy in VQA (52.59 vs. 51.84). Thus, even with a dataset consisting of only 2 shots, ICD-LM is still able to generate longer high-quality ICD configurations. These experiment results indicate that ICD-LM can effectively generate optimal ICD configurations, thereby assisting LVLM in predictions.
Furthermore, we observe the ICD configurations generated by ICD-LM are more robust. We can observe the SITR underperforms when compared to RS in both IC and VQA.
For STTR in VQA, it exhibits a performance decline compared to RS when the number of shots is relatively large. For instance, it lags behind RS by 0.81 in the 8-shot configuration. These results highlight that not all similarity-based retrieval methods are effective in different multimodal tasks.
In Fig. 4, we visualize several ICD configurations and observe that, when the context of the ICDs is similar to query sample, OpenFlamingo tends to excessively rely on text information of ICDs when predicting labels. For example, in Fig. 4 (a) and (b), all the ICD questions are yes-or-no type, which causes LVLMs to output “no” even when faced with a “What” type question of the query sample. While for IC in Fig. 4 (c) and (d), SITR incorrectly outputs “London” even though the location is not explicitly indicated in the query image, while SIIR even leads LVLMs to directly copy the text from the ICD. This behavior, identified as Short-cut Inference, has been previously highlighted in a study [41]. Conversely, ICD-LM generates more diverse ICD configurations, thereby preventing misleading inferences.
Table 2: Results of diverse ICL methods on IC and VQA.
Interpolation Extrapolation Avg:1$\sim$8
Shot 1 Shot 2 Avg:1$\sim$2 Shot 3 Shot 4 Shot 6 Shot 8 Avg:3$\sim$8
IC RS 73.32 82.95 78.14 87.72 93.65 95.81 97.42 93.65 88.48
SITR 66.05 77.69 71.87 83.46 85.05 89.84 93.57 87.98 82.61
SIIR 71.71 84.87 78.29 90.83 93.22 97.80 101.01 95.72 89.91
ICD-LM 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
VQA RS 41.97 45.92 43.95 48.17 48.95 51.18 51.44 49.94 47.94
SITR 40.17 43.58 41.88 46.03 47.5 49.72 50.75 48.50 46.29
SIIR 43.31 47.46 45.39 49.85 50.68 53.23 53.58 51.84 49.69
STTR 44.6 46.75 45.68 47.92 49.05 50.06 49.47 49.13 47.98
ICD-LM 46.66 50.83 48.75 51.91 52.15 53.29 53.01 52.59 51.31
4.4 Ablation Studies
We use ablation studies to explore the effects of diverse settings on our approach, including (1) diverse $\mathcal{D}_{\mathcal{M}}$ configurations; (2) diverse scorers $I_{\mathcal{M}}$ in Eq. (3); (3) diverse LM structures; (4) $\mathcal{D}_{\mathcal{M}}$ with longer few-shot ICDs; and (5) randomly ordering the ICD-LM generated ICD sequences.
Diverse $\mathcal{D}_{\mathcal{M}}$ Configurations. We generate 2-shot $\mathcal{D}_{\mathcal{M}}$ by different settings to investigate the corresponding effects. We select three factors for our ablation studies: beam size $b$; the number $n$ of samples in $\mathcal{A}$; and the sampling method of $\mathcal{D}_{S}^{\bm{a}}$, which includes 3 methods: Random: Selecting randomly from $\mathcal{D}_{S}$; Similar Text (Sim-T): Selecting the highest textual similarity sample with anchor sample $\bm{a}$ from $\mathcal{D}_{S}$; Similar Image (Sim-I): Selecting the highest visual similarity sample with anchor sample $\bm{a}$ from $\mathcal{D}_{S}$.
Table 34 show the results for different $\mathcal{D}_{\mathcal{M}}$ configurations on IC and VQA, respectively. We find that ICD-LM can consistently improve the performance compared with the baseline RS The RS means Random Sample ICD retrieval method which is mentioned in Section 4.3.1. in Avg:1$\sim$2. As for length extrapolation capability, only Sim-T gets a lower score than RS. These comparisons confirm the robustness of our method.
Besides, we observe that increasing the beam size has a positive correlation with ICD-LM performance. Specifically, as the beam size increases from $1$ to $5$, the CIDEr score increases by 2.79 for IC and the accuracy increases 1.73 in VQA in Avg:1$\sim$8. This suggests that a diverse $\mathcal{D}_{\mathcal{M}}$ encompasses a broader range of high-quality ICD configurations, which can help train a better ICD-LM. However, an excessively large beam size can negatively impact performance. For instance, in VQA, the accuracy of $b=10$ decays 0.12 than $b=5$ in Avg:1$\sim$8. We hypothesize this drop in performance is due to the introduction of lower-scoring ICD sequences with a large beam size, potentially misleading ICD-LM during training.
Moreover, we observe that using more anchor samples can improve the interpolation performance in both IC and VQA, e.g., when $n$ increases from 1000 to 5000, the Avg:1$\sim$2 CIDEr/accuracy of IC/VQA increases from 83.94/45.39 to 84.32/ 48.75. However, we find that on IC, although the interpolation performance increases when $n$ changes from 3000 to 5000, the extrapolation performance decays, e.g., Avg:3$\sim$8 decrease from 97.60 to 96.52. One possible reason is that using Eq. (3) to build $D_{\mathcal{M}}$ may introduce certain in-domain bias which is beneficial for interpolation while detrimental for extrapolation on IC. Then using more such data may enhance such in-domain bias and thus damage the extrapolation.
For different sample methods of constructing the $\mathcal{D}_{S}^{a}$, we find Random is the best in both IC and VQA. We suppose this is because selecting similar ICDs with the anchor sample from $\mathcal{D}_{S}$ will damage the diversity. Previous study [12] in NLP validates that the diversity of the ICD sequences will also help improve the performance of LLMs.
Table 3: CIDEr of diverse $\mathcal{D}_{\mathcal{M}}$ configurations on IC.
Interpolation Extrapolation Avg:1$\sim$8
Shot 1 Shot 2 Avg:1$\sim$2 Shot 3 Shot 4 Shot 6 Shot 8 Avg:3$\sim$8
RS 73.32 82.95 78.1 87.72 93.65 95.81 97.42 93.6 88.5
$b=1$ 75.67 84.15 79.91 90.10 92.93 95.92 99.16 94.53 89.66
$b=5$ 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
$b=10$ 80.64 89.27 84.96 94.47 96.26 98.58 99.15 97.12 93.06
$n=1000$ 79.55 88.34 83.94 91.16 96.25 99.70 99.87 96.74 92.48
$n=3000$ 79.30 88.96 84.13 93.28 97.66 99.34 100.15 97.60 93.11
$n=5000$ 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
Sim-I 75.32 88.61 81.96 93.28 95.28 97.71 98.19 96.11 91.40
Sim-T 75.78 86.66 81.22 84.39 84.27 88.47 93.53 87.66 85.52
Random 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
Table 4: Accuracy of diverse $\mathcal{D}_{\mathcal{M}}$ configurations on VQA.
Interpolation Extrapolation Avg:1$\sim$8
Shot 1 Shot 2 Avg:1$\sim$2 Shot 3 Shot 4 Shot 6 Shot 8 Avg:3$\sim$8
RS 41.97 45.92 43.95 48.17 48.95 51.18 51.44 49.94 47.94
$b=1$ 44.56 48.38 46.47 49.76 49.95 52.27 52.54 51.13 49.58
$b=5$ 46.66 50.83 48.75 51.91 52.15 53.29 53.01 52.59 51.31
$b=10$ 46.89 50.27 48.58 51.54 52.17 53.00 53.26 52.49 51.19
$n=1000$ 44.66 46.12 45.39 47.96 50.15 52.09 51.56 50.44 48.76
$n=3000$ 46.04 49.08 47.56 50.42 50.88 51.54 51.59 51.11 49.93
$n=5000$ 46.66 50.83 48.75 51.91 52.15 53.29 53.01 52.59 51.31
Sim-I 45.00 49.19 47.10 50.49 51.38 52.64 52.66 51.79 50.23
Sim-T 44.30 46.45 45.38 48.36 48.95 50.46 50.44 49.55 48.16
Random 46.66 50.83 48.75 51.91 52.15 53.29 53.01 52.59 51.31
Diverse Scorers $I_{\mathcal{M}}$ for evaluating ICD sequences. To evaluate the quality of ICD configurations, we can also utilize task-specific metrics as $I_{\mathcal{M}}$ to build $D_{\mathcal{M}}$, such as the CIDEr metric in IC. Table 5 compares the results obtained by using prediction confidence and CIDEr as $I_{\mathcal{M}}$.
Table 5: CIDEr of diverse scorers on IC.
Interpolation Extrapolation Avg:1$\sim$8
Shot 1 Shot 2 Avg:1$\sim$2 Shot 3 Shot 4 Shot 6 Shot 8 Avg:3$\sim$8
RS 73.32 82.95 78.14 87.72 93.65 95.81 97.42 93.65 88.48
Confidence 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
CIDEr 84.86 90.99 87.93 92.53 94.25 94.98 92.31 93.52 91.65
We can observe that the ICD-LM trained with $\mathcal{D}_{\mathcal{M}}$ using CIDEr scorer can achieve 3.61 higher than Confidence scorer in Avg:1$\sim$2, which means that the CIDEr scorer can assign a more accurate and reasonable score for ICD configurations. However, the length extrapolation capability decreases obviously, which is 3.0 lower than Confidence score in Avg:3$\sim$8. These results evidently show the robustness of Confidence scorer. Although the result shows that the task-specific metric can help the ICD-LM to improve the interpolation ability, it will cost more time and more GPU memory to construct $\mathcal{D}_{\mathcal{M}}$. Specifically, in IC, the time cost of constructing $\mathcal{D}_{\mathcal{M}}$ is approximately 10 times of Confidence.
Diverse LM Structures. In Table  6, we use LSTM [27] as the structure of ICD-LM to generate ICDs for IC and VQA, respectively. We can find that even when using LSTM, our method still achieves excellent performance. In IC, the overall performance improves by 3.64 (92.12 vs. 88.48) compared to the RS baseline, while in VQA, it is improved by 1.38 (49.32 vs. 47.94). However, due to the weak representation learning capability of LSTM, its performance is lower than the Transformer structure,e.g., the scores decrease by 0.33 and 1.99 in IC and VQA, respectively. Overall, these results demonstrate the ability to generate high-quality ICD configurations using various LMs. This validates that employing LM for ICD configuration is a reasonable and effective approach.
Table 6: Results of diverse LM structures on IC and VQA.
Interpolation Extrapolation Avg:1$\sim$8
Shot 1 Shot 2 Avg:1$\sim$2 Shot 3 Shot 4 Shot 6 Shot 8 Avg:3$\sim$8
IC RS 73.32 82.95 78.14 87.72 93.65 95.81 97.42 93.65 88.48
ICD-LM(LSTM) 79.73 88.14 83.93 92.36 96.89 96.99 98.63 96.21 92.12
ICD-LM(Transformer) 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
VQA RS 41.97 45.92 43.95 48.17 48.95 51.18 51.44 49.94 47.94
ICD-LM(LSTM) 44.64 48.55 46.60 49.77 50.51 50.71 51.72 50.68 49.32
ICD-LM(Transformer) 46.66 50.83 48.75 51.91 52.15 53.29 53.01 52.59 51.31
Table 7: Results of ICD-LM with 4-shot $\mathcal{D}_{\mathcal{M}}$ on IC and VQA.
Interpolation Extrapolation Avg:1$\sim$8
Shot 1 Shot 2 Shot 3 Shot 4 Avg:1$\sim$4 Shot 6 Shot 8 Avg:6$\sim$8
IC RS 73.32 82.95 87.72 93.65 84.41 95.81 97.42 96.62 88.48
SITR 66.05 78.66 85.47 86.47 79.16 92.27 95.63 93.95 84.09
SIIR 71.32 84.03 91.38 94.28 85.25 98.63 101.89 100.26 90.25
ICD-LM(4-shot $\mathcal{D}_{\mathcal{M}}$) 75.73 85.97 91.28 96.41 87.35 97.56 98.35 97.96 90.88
ICD-LM(2-shot $\mathcal{D}_{\mathcal{M}}$) 80.02 88.63 93.41 96.06 89.53 97.26 99.35 98.30 92.45
VQA RS 41.97 45.92 48.17 48.95 46.25 51.18 51.44 51.31 47.94
SITR 39.84 43.60 46.12 47.34 44.23 49.54 50.49 50.02 46.16
SIIR 43.05 47.25 49.68 50.63 47.65 53.16 53.28 53.22 49.51
STTR 43.53 46.89 48.13 48.87 46.86 49.23 49.97 49.60 47.77
ICD-LM(4-shot $\mathcal{D}_{\mathcal{M}}$) 44.22 48.11 49.30 51.12 48.19 52.17 52.76 52.47 49.61
ICD-LM(2-shot $\mathcal{D}_{\mathcal{M}}$) 46.66 50.83 51.91 52.15 50.39 53.29 53.01 53.15 51.31
Table 8: Results of Random Order of ICD-LM generated ICDs.
2-Shot $\mathcal{D}_{\mathcal{M}}$ 4-Shot $\mathcal{D}_{\mathcal{M}}$
VQA IC VQA IC
Original 50.83 88.63 51.12 85.97
Random Order 50.42 88.56 50.63 85.77
Table 9: Results of Fixed Set ICD-LM on IC and VQA.
Interpolation Extrapolation Avg:1$\sim$8
Shot 1 Shot 2 Avg:1$\sim$2 Shot 3 Shot 4 Shot 6 Shot 8 Avg:3$\sim$8
IC RS 73.32 82.95 78.14 87.72 93.65 95.81 97.42 93.65 88.48
SIIR 71.71 84.87 78.29 90.83 93.22 97.80 101.01 95.72 89.91
non Fixed Set 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
Fixed Set-1 75.41 88.14 81.78 92.48 94.20 99.22 103.86 97.44 92.22
Fixed Set-2 86.37 96.03 91.20 89.80 99.81 105.39 103.51 99.63 96.82
VQA RS 41.97 45.92 43.95 48.17 48.95 51.18 51.44 49.94 47.94
SIIR 43.31 47.46 45.39 49.85 50.68 53.23 53.58 51.84 49.69
non Fixed Set 46.66 50.83 48.75 51.91 52.15 53.29 53.01 52.59 51.31
Fixed Set-1 44.24 51.32 47.78 50.20 51.60 54.39 53.83 52.51 50.93
Fixed Set-2 44.24 46.40 45.32 48.38 48.41 49.16 50.23 49.05 47.80
Figure 4: Visualizations of using ICDs configured by different methods on VQA. Due to the limited space, we only show the first and the last ICDs. We can find that ICD-LM tends to use more diverse ICDs and thus not lead to short-cut inference.
Longer Few-shot $\mathcal{D}_{\mathcal{M}}$. We further explore the performance of ICD-LM using a 4-shot $\mathcal{D}_{\mathcal{M}}$, as presented in Table 7. It is evident that ICD-LM continues to outperform other retrieval-based methods in Avg:1$\sim$8 metric.
However, we observe a notable performance reduction when the ICD-LM is trained with the 4-shot $\mathcal{D}_{\mathcal{M}}$ compared to the 2-shot $\mathcal{D}_{\mathcal{M}}$. Specifically, for IC, there is a performance decrease of approximately 1.57 in the Avg:1$\sim$8 metric when using the 4-shot $\mathcal{D}_{\mathcal{M}}$ to train ICD-LM. We speculate that this is due to the use of Eq. (2), which may lead to becoming stuck in a local optimum when constructing longer few-shot sequences. Specifically, when selecting the $k$-th ICD, the previous $k-1$ ICDs are held constant. This selection process might not always result in the optimal $k$-shot sequence. When shot number increases, the ICD sequence may further deviate from the global optimum, leading to a decline in ICD-LM performance. Thus leading to poor performance of the ICD-LM generated by this dataset.
Random Order ICD sequence. To validate that our model achieves an effective ICD order, we randomly rearrange the ICD sequences generated by ICD-LM and evaluate the performance. Specifically, we use the ICD-LM trained with 2-shot and 4-shot $\mathcal{D}_{\mathcal{M}}$ and evaluate their performance of the 2-shot and 4-shot ICD configurations, respectively. Table 8 displays our results. It is evident that the original order of ICD configurations generated by the ICD-LM attains the highest score in both VQA and IC, demonstrating that ICD-LM can effectively learn how to order the ICDs.
Fixed Set. During training ICD-LM, we discover that using a high learning rate and not freezing the CLIP model may make ICD-LM converge to a specific solution that for any query input, the ICD configuration is the same and we term this case “Fixed Set”. Table 9 presents the performance of multiple Fixed Set models obtained with different training parameters. The specific training parameter configurations and the generated ICD configurations will be shown in Appendix. We find that the performance of the Fixed Set models fluctuates significantly. For example, in IC, the best version, Fixed Set-2, can outperform the non-Fixed Set by 4.6 points in Avg:1$\sim$8. However, the performance of Fixed Set-2 is poorer than RS in VQA. This phenomenon of Fixed Set demonstrates that for a specific LVLM, there exists a particular ICD configuration that can universally enhance ICL capability. Therefore, we do not need to store a large number of samples for ICD retrieval; instead, we only need to use the Fixed Set samples as the core supporting set. Therefore, it is worthwhile to consider further research in the future to stabilize the performance of the Fixed Set.
5 Conclusion and Limitations
In this paper, motivated from the observation that configuring an ICD sequence is a mirror process of composing a sentence, we propose to build an ICD-LM for configuring ICDs. To achieve this, we construct a dataset containing high-quality ICD sequences to train this ICD-LM. After training, we validate the effectiveness of ICD-LM in accomplishing ICD selection and ordering by comparing it with similarity-based retrieval methods. Also, we deploy extensive ablations to validate the effectiveness of diverse settings of dataset construction and model building strategies.
One major limitation of our study is the strategy used to build $\mathcal{D}_{\mathcal{M}}$ is not optimal, which requires further improvement. This limitation is revealed by observing that the 4-shot $\mathcal{D}_{\mathcal{M}}$ performs worse than the 2-shot one, highlighting the need for a more effective approach in searching for longer ICD sequences. To address this, we plan to design function $I_{\mathcal{M}}$ to evaluate the effectiveness of ICD sequences; and use better sampling strategies in Eq. (2) to avoid the ICD sequence deviating from the global optimum.
Supplementary Material
Implementation Details. We provide additional implementation details for each experiment in Table.  10. For all experiments, the batch size is set to 64, the warmup steps to $5\%$ of total training steps, the scheduler as a cosine scheduler, and the optimizer as AdamW [19].
Table 10: Different settings of ICD-LM experiments, where the $n$ is the number of anchor samples in $\mathcal{A}$, $b$ is the beam size, and $l$ is the length of ICD configurations.
Training Parameters $\mathcal{D}_{\mathcal{M}}$ Parameters Avg:1$\sim$8
lr epoch Freeze Adapter $n$ $b$ $l$
[…]
Fixed Set ICD Configurations. We present four ICD configurations of the Fixed Set. Figure 5 displays two ICD configurations for IC, and Figure 6 displays the other two ICD configurations for VQA.
Figure 5: 8-shot ICD configurations Visualizations of IC Fixed Set.
Figure 6: 8-shot ICD configurations Visualizations of VQA Fixed Set.